# Reliability, and Global Privacy

### v3

@@ X Conclusion
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\left(e^{\frac{c_{5}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J},\mathcal{E})+\delta_{f^{\prime},j}^{(5)}\right)$ (93a)
$\leq e^{\frac{c_{5}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J})+(\delta_{f^{\prime},j}^{(5)}+\delta^{\prime})$ (93b)
@@
$\leq e^{\frac{c_{5}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J}^{c})$ (94)
@@
$\leq e^{\frac{c_{5}}{\sqrt{\bar{\nu}_{f^{\prime}}^{t}-r_{2}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime})+\mathrm{P_{r}}(\mathcal{J})(\delta_{f^{\prime},j}^{(5)}+\delta^{\prime}),$ (95)
@@
Appendix E Proof of Theorem VI.5
@@
When $\Delta\bm{\Theta}_{j}^{t}$ is perturbed from $\bm{u}$ to $\bm{u}^{\prime}$, the the $\ell_{2}$-sensitivity can be bounded by
$\left\lVert\frac{1}{K}(\bm{u}-\bm{u}^{\prime})\right\rVert\leq\frac{2\zeta}{K}\sqrt{D(1+\delta_{0})},$ (98)
with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$. So $\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ is $(\epsilon^{(7)},\delta^{(7)})$-differentially private in protecting the local data of client $j$ with probability $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where
@@ Appendix F Choices of Bernstein Parameters
F-A Choice of $r_{1}$
Let $X_{k,1}=\nu_{k}^{t}-\bar{\nu}_{k}$. Then, $\mathbb{E}[X_{k,1}]=0$, and $\max\{X_{k,1}\}\leq\max\{\nu_{k}^{t}\}\leq\max_{k}\{\nu_{k}^{t}\}\triangleq M_{1}$. Due to Bernoulli r.v.s, $M_{1}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,1}^{2}]=\mathbb{E}[(\nu_{k}^{t}-\bar{\nu}_{k})^{2}]=\mathbb{E}[(\nu_{k}^{t})^{2}]-\bar{\nu}_{k}^{2},$ (103)
where $\bar{\nu}_{k}$ and $\mathbb{E}[(\nu_{k}^{t})^{2}]$ are given in (122) and (125b), respectively. Thereby, $r_{1}$ is chosen as
@@ F-B Choice of $r_{2}$
Let $X_{k,2}=\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}}$. Then, $\mathbb{E}[X_{k,2}]=0$, and $\max\{X_{k,2}\}\leq\max\{\nu_{f^{\prime}}^{t}\}\leq\max_{f^{\prime}}\{\nu_{f^{\prime}}^{t}\}\triangleq M_{2}$. Due to Bernoulli r.v.s, $M_{2}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,2}^{2}]=\mathbb{E}[(\nu_{f^{\prime}}^{t}-\bar{\nu}_{f^{\prime}})^{2}]=\mathbb{E}[(\nu_{f^{\prime}}^{t})^{2}]-\bar{\nu}_{f^{\prime}}^{2},$ (105)
@@
where $\bar{\nu}_{f^{\prime}}$ and $\mathbb{E}[(\nu_{f^{\prime}}^{t})^{2}]$ are given in (129) and (125b), respectively. Thereby, $r_{2}$ is chosen as
F-C Choice of $r_{3}$
According to Appendix E, we have $r_{3}=\zeta\sqrt{D(1+\delta_{0})}$.
Appendix G Proof of Theorem VIII.1
Under Assumption VIII.1, it holds that
Next, we bound the two terms in (107), respectively.
G-A Bounding the First Term
For the first term, we have
$=\mathbb{E}\hskip-2.84526pt\left[\hskip-2.84526pt\left\langle\hskip-2.84526pt\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),-\eta\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{R_{t_{s}}I}a_{i}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1},\bm{\xi}_{k}^{t_{s},i})\hskip-2.84526pt\right\rangle\hskip-2.84526pt\right]$
$=-\eta\hskip-2.84526pt\sum_{i=1}^{R_{t_{s}}I}\hskip-2.84526pta_{i}\mathbb{E}\hskip-2.84526pt\left[\hskip-2.84526pt\left\langle\hskip-2.84526pt\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\frac{1}{K}\hskip-2.84526pt\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1})\hskip-2.84526pt\right\rangle\hskip-2.84526pt\right]$ (108a)
$\leq-\frac{1}{2}\eta\sum_{i=1}^{R_{t_{s}}I}a_{i}\mathbb{E}\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt+\frac{1}{2}\eta\hskip-2.84526pt\sum_{i=1}^{R_{t_{s}}I}a_{i}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\hskip-2.84526pt\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}_{k}^{t_{s}-1})\hskip-2.84526pt-\hskip-2.84526pt\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s},i-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt-\frac{1}{2}\eta\sum_{i=1}^{R_{t_{s}}I}a_{i}\mathbb{E}\left[\left\lVert\frac{1}{K}\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s},i-1})\right\lVert^{2}\right]$ (108b)
$\leq-\frac{1}{2}\eta\sum_{i=1}^{R_{t_{s}}I}a_{i}\mathbb{E}\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt+\frac{1}{2}\eta\frac{G^{2}}{K}\sum_{i=1}^{R_{t_{s}}I}a_{i}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert\bm{\Theta}_{k}^{t_{s}-1}-\bm{\Theta}^{t_{s},i-1}\right\lVert^{2}\right],$ (108c)
where (108a) follows Assumption VIII.2, (108b) is due to the convexity of $\ell_{2}$-norm, and (108c) is due to Assumption VIII.1.
G-B Bounding the Second Term
For the second term, it holds that
$=\frac{G}{2}\mathbb{E}\left[\left\lVert-\eta\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{R_{t_{s}}I}a_{i}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1},\bm{\xi}_{k}^{t,i})\right\rVert^{2}\right]$
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert\sum_{i=1}^{R_{t_{s}}I}a_{i}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1},\bm{\xi}_{k}^{t_{s},i})\right\rVert^{2}\right]$ (109a)
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i=1}^{R_{t_{s}}I}a_{i}\mathbb{E}\hskip-1.42262pt\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1},\bm{\xi}_{k}^{t_{s},i})\right\rVert^{2}\right]$ (109b)
$=\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i=1}^{R_{t_{s}}I}a_{i}\Bigg{(}\mathbb{E}\bigg{[}\bigg{\lVert}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1},\bm{\xi}_{k}^{t_{s},i})$
$\hskip 14.22636pt-\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1})\bigg{\rVert}^{2}\bigg{]}+\mathbb{E}\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1})\right\rVert^{2}\right]\Bigg{)}$ (109c)
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i=1}^{R_{t_{s}}I}a_{i}\Bigg{(}\sigma^{2}+\mathbb{E}\bigg{[}\bigg{\lVert}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i-1})$
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}+\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i=1}^{R_{t_{s}}I}a_{i}\cdot$
$\bigg{(}G^{2}\mathbb{E}\bigg{[}\bigg{\lVert}\bm{\Theta}_{k}^{t_{s},i-1}-\bm{\Theta}^{t_{s}-1}\bigg{\rVert}^{2}\bigg{]}+\mathbb{E}\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s}-1})\right\rVert^{2}\right]\Bigg{)}$ (109e)
$+\frac{\eta^{2}G^{3}}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i=1}^{R_{t_{s}}I}a_{i}\mathbb{E}\bigg{[}\bigg{\lVert}\bm{\Theta}_{k}^{t_{s},i-1}-\bm{\Theta}^{t_{s}-1}\bigg{\rVert}^{2}\bigg{]},$ (109f)
where (109a) and (109b) use Jensen’s Inequality, (109c) is due to unbiased estimation, (109d) follows Assumption VIII.2, (109e) follows Assumption VIII.1, and (109f) follows Assumption VIII.3.
G-C Immediate Result
Substitute (108c) and (109f) into (107), then substitute Lemma VIII.1, we obtain the immediate result for one round,
Using Assumption VIII.3, this further becomes
G-D Average over Successful Rounds
G-E Relaxation with Probability Guarantee
By (120), (113) can be further bounded as follows with probability $0.9974$,
Substitute $e_{R},\nu_{R},e_{R^{2}},\nu_{R^{2}}$, we get
$\frac{1}{T_{s}}\hskip-2.84526pt\sum_{t_{s}=1}^{T_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t-1})\rVert^{2}\right]\hskip-1.42262pt\leq\hskip-1.42262pt\frac{1}{T_{s}}\hskip-2.84526pt\sum_{t_{s}=1}^{T_{s}}\hskip-2.84526ptR_{t_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t-1})\rVert^{2}\right]$
Appendix H Renewal Theory
According to the algorithm’s stopping rule, training terminates at the first round in which the global model is successful and the cumulative number of training rounds reaches the threshold $T$. Hence, we have
$T_{s}=\min\left\{n:\sum_{t=1}^{n}R_{t_{s}}\geq T\right\},$ (116)
$e_{R^{2}}=\mathbb{E}[R_{t}^{2}]=\hskip-5.69054pt\sum_{R_{t}=1}^{+\infty}\hskip-2.84526ptR_{t}^{2}P_{O}^{R_{t}-1}(1-P_{O})=\hskip-2.84526pt\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-2}(P_{O}),$ (117a)
$\nu_{R^{2}}=\mathrm{var}(R_{t}^{2})=\mathbb{E}[R_{t}^{4}]-e_{R^{2}}^{2}$
According to the strong law of large numbers, the empirical average of any function $f(R)$ over the first $T_{s}$ renewals converges to its expected value almost surely, i.e., $\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}\rightarrow\mathbb{E}[R_{t}^{2}]$. According to CLT,
Appendix I Calculation of Privacy Noise Variance
I-A Calculation of $\bar{\nu}_{k}$ and $\mathbb{E}[(\nu_{k}^{t})^{2}]$
@@
Let us begin by computing the variance of the aggregated secret keys $\bm{M}_{k}^{t}$ at relaying client $k$. It can be verified that $\bm{M}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}=\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}$, so $\mathrm{var}(\bm{M}_{k}^{t})=\nu_{k}^{t}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\bm{I}_{D}$ in (34). Similarly, we define $\bm{M}_{k}^{t,-j}=\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}\bm{N}_{m}^{t}=\sum_{m\neq j}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}$, and we have $\mathrm{var}(\bm{M}_{k}^{t,-j})=\nu_{k}^{t,-j}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\bm{I}_{D}$ in (LABEL:eq:_agg_noise_no_j). The $\bm{\Lambda}_{k}^{t}$ depends on $\{\tau_{k,m}\}_{m=1}^{K}$, so the variance $\nu_{k}^{t}$ of the effective noise $\bm{M}_{k}^{t}$ is also a random variable concerning the network topology. It should be noted that due to correlation, the variance of $\bm{M}_{k}^{t}$ does not necessarily increase monotonously with the number of received updates $|\mathcal{U}_{k}^{t}\rvert$ at client $k$. The expected variance $\bar{\nu}_{k}$ is given by
@@
I-B Calculation of $\bar{\nu}_{f^{\prime}}$ and $\mathbb{E}[(\nu_{f^{\prime}}^{t})^{2}]$

### v4

@@ VIII Conclusion
$\leq\delta^{\prime}+\mathrm{P_{r}}(\mathcal{E})\left(e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t}-r_{4}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J},\mathcal{E})+\delta_{f^{\prime},j}^{(5)}\right)$ (81a)
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t}-r_{4}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J})+(\delta_{f^{\prime},j}^{(5)}+\delta^{\prime})$ (81b)
@@
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t}-r_{4}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime},\mathcal{J}^{c})$ (82)
@@
$\leq e^{\frac{c_{4}}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t}-r_{4}}}}\mathrm{P_{r}}(\bm{F}_{f^{\prime}}^{t}\hskip-2.84526pt\in\hskip-2.84526pt\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}\hskip-2.84526pt=\bm{u}^{\prime})+\mathrm{P_{r}}(\mathcal{J})(\delta_{f^{\prime},j}^{(5)}+\delta^{\prime}),$ (83)
@@
Appendix E Proof of Theorem V.5
@@
The (84) holds before and after perturbation. Hence, $\mathrm{P_{r}}\left(\bm{u}\in\mathbb{B}_{D}(\zeta\sqrt{D(1+\delta_{0})})\right)\geq 1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ and $\mathrm{P_{r}}\left(\bm{u^{\prime}}\in\mathbb{B}_{D}(\zeta\sqrt{D(1+\delta_{0})})\right)\geq 1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$. When $\Delta\bm{\Theta}_{j}^{t}$ is perturbed from $\bm{u}$ to $\bm{u}^{\prime}$, the $\ell_{2}$-sensitivity can be bounded by
$\left\lVert\frac{1}{K}(\bm{u}-\bm{u}^{\prime})\right\rVert\leq\frac{2\zeta}{K}\sqrt{D(1+\delta_{0})}.$ (86)
with probability $\left(1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}\right)^{2}$. So $\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ is $(\epsilon^{(7)},\delta^{(7)})$-differentially private in protecting the local data of client $j$ with probability $\left(1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}\right)^{2}$, where
@@ Appendix F Choices of Bernstein Parameters
F-A Choice of $r_{1}$ and $r_{2}$
Let $X_{k,2}=\nu_{k}^{t}-\bar{\nu}_{k}$. Then, $\mathbb{E}[X_{k,2}]=0$, and $\max\{X_{k,2}\}\leq\max\{\nu_{k}^{t}\}\leq\max_{k}\{\nu_{k}^{t}\}\triangleq M_{2}$. Due to Bernoulli r.v.s, $M_{2}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,2}^{2}]=\mathbb{E}[(\nu_{k}^{t}-\bar{\nu}_{k})^{2}]=\mathbb{E}[(\nu_{k}^{t})^{2}]-\bar{\nu}_{k}^{2},$ (91)
where $\bar{\nu}_{k}$ and $\mathbb{E}[(\nu_{k}^{t})^{2}]$ are given in (100) and (103b), respectively. Thereby, $r_{2}$ is chosen as
Similarly, let $X_{k,1}=\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}$. Then, $\mathbb{E}[X_{k,1}]=0$, and $\max\{X_{k,1}\}\leq\max\{\nu_{k}^{t,-j}\}\leq\max_{k}\{\nu_{k}^{t,-j}\}\triangleq M_{1}$. By setting $\tau_{k,j}=0$, $M_{1}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,1}^{2}]=\mathbb{E}[(\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j})^{2}]=\mathbb{E}[(\nu_{k}^{t,-j})^{2}]-(\bar{\nu}_{k}^{-j})^{2},$ (93)
where $(\bar{\nu}_{k}^{-j})^{2}$ is given in (102) and $\mathbb{E}[(\nu_{k}^{t,-j})^{2}]$ can be computed by setting $\tau_{k,j}=0$ in (103b). Thereby, $r_{1}$ is chosen as
@@ F-B Choice of $r_{3}$ and $r_{4}$
Let $X_{k,4}=\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}$. Then, $\mathbb{E}[X_{k,4}]=0$, and $\max\{X_{k,4}\}\leq\max\{\nu_{F_{f^{\prime}}}^{t}\}\leq\max_{f^{\prime}}\{\nu_{F_{f^{\prime}}}^{t}\}\triangleq M_{4}$. Due to the Bernoulli r.v.s, $M_{4}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,4}^{2}]=\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}})^{2}]=\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t})^{2}]-\bar{\nu}_{F_{f^{\prime}}}^{2},$ (95)
where $\bar{\nu}_{F_{f^{\prime}}}$ and $\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t})^{2}]$ are given in (107) and (103b), respectively. Thereby, $r_{4}$ is chosen as
$r_{4}=\ln\left(\frac{2}{\delta^{\prime}}\right)\cdot\left(\frac{1}{3}M_{4}+\sqrt{\frac{1}{9}M_{4}^{2}+\frac{\mathbb{E}[X_{k,4}^{2}]}{\ln\left(\frac{2}{\delta^{\prime}}\right)}}\right).$ (96)
Similarly, let $X_{k,3}=\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}}$. Then, $\mathbb{E}[X_{k,3}]=0$, and $\max\{X_{k,3}\}\leq\max\{\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}\}\leq\max_{f^{\prime}}\{\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}\}\triangleq M_{3}$. By setting $\tau_{k,j}=0$, $M_{3}$ is computable, through e.g., brute-force search. Moreover,
$\mathbb{E}[X_{k,3}^{2}]$ $=\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}})^{2}]$
$=\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t,\{k,-j\}})^{2}]-(\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}})^{2},$ (97)
@@
where $\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}}$ and $\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t,\{k,-j\}})^{2}]$ can be computed be setting $\tau_{k,j}=0$ in (107) and (112), respectively. Thereby, $r_{3}$ is chosen as
$r_{3}=\ln\left(\frac{2}{\delta^{\prime}}\right)\cdot\left(\frac{1}{3}M_{3}+\sqrt{\frac{1}{9}M_{3}^{2}+\frac{\mathbb{E}[X_{k,3}^{2}]}{\ln\left(\frac{2}{\delta^{\prime}}\right)}}\right).$ (98)
Appendix G Calculation of Privacy Noise Variance
G-A Calculation of $\bar{\nu}_{k}$ and $\mathbb{E}[(\nu_{k}^{t})^{2}]$
@@
Let us begin by computing the variance of the aggregated secret keys $\bm{M}_{k}^{t}$ at relaying client $k$. It can be verified that $\bm{M}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}=\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}$, so $\mathrm{var}(\bm{M}_{k}^{t})=\nu_{k}^{t}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\bm{I}_{D}$ in (LABEL:eq:_agg_noise). Similarly, we define $\bm{M}_{k}^{t,-j}=\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}\bm{N}_{m}^{t}=\sum_{m\neq j}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}$, and we have $\mathrm{var}(\bm{M}_{k}^{t,-j})=\nu_{k}^{t,-j}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\bm{I}_{D}$ in (LABEL:eq:_agg_noise_no_j). The $\bm{\Lambda}_{k}^{t}$ depends on $\{\tau_{k,m}\}_{m=1}^{K}$, so the variance $\nu_{k}^{t}$ of the effective noise $\bm{M}_{k}^{t}$ is also a random variable concerning the network topology. It should be noted that due to correlation, the variance of $\bm{M}_{k}^{t}$ does not necessarily increase monotonously with the number of received updates $|\mathcal{U}_{k}^{t}\rvert$ at client $k$. The expected variance $\bar{\nu}_{k}$ is given by
@@
G-B Calculation of $\bar{\nu}_{F_{f^{\prime}}}$ and $\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t})^{2}]$